OpenStudy (anonymous):

Factor by grouping. 7x(5x-5)-1(7x-1)

OpenStudy (anonymous):

is it...35x^2 - 42x + 1??!!

OpenStudy (anonymous):

but, you're supposed to factor it not simplify aren't you ?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

(7x-1)(7x-1)......no?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=35x^2-42x-1 That has ugly roots ...

OpenStudy (anonymous):

are you sure it's written correctly ?

OpenStudy (anonymous):

hmm 35x^2-32x+5=

OpenStudy (anonymous):

I mean is the original question written correctly ? See the ugly roots we got for it ? http://www.wolframalpha.com/input/?i=7x%285x-5%29-1%287x-1%29

OpenStudy (anonymous):

original question is that problem.

OpenStudy (anonymous):

thanks for trying, i must of had goofed somewhere in the problem.

OpenStudy (anonymous):

but homework is past due now so oh well...XD maybe i'll figure it out later.

OpenStudy (anonymous):

was that question from a text book ? or a site ?

OpenStudy (anonymous):

math program

OpenStudy (anonymous):

there were three steps to it I had to answer..I think I messed up on the question factor part

OpenStudy (anonymous):

but this question came from that program ? factor by grouping: 7x(5x-5)-1(7x-1)

OpenStudy (anonymous):

no that's the problem that I got from trying to factor by grouping.

OpenStudy (anonymous):

original problem is the 35 one

OpenStudy (anonymous):

This is the original ? 35x^2 - 42x + 1

OpenStudy (anonymous):

35x^2-32x+5

OpenStudy (anonymous):

that comes out weird too :(

OpenStudy (anonymous):

odd....maybe it's a mistake from the program. let me double check what the problem is

OpenStudy (anonymous):

yeah it told me to complete the parts below to factor that problem by grouping, first part was finding the two numbers who's product is 35.3=173 and whose sum is -32. I got -25, and -7, and then it ask me to write -32x using the factors so it was -25x-7x, and then it told me to factor by grouping, and dats when I messed up.

OpenStudy (anonymous):

(-25)*(-7) = 135

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

isnt it : x^2-32x+135 ?

OpenStudy (anonymous):

i have no idea..my brain is fried. XD I got everything write but the factoring part.

OpenStudy (anonymous):

x^2-32x+135 = (x-5)(x-27)

OpenStudy (anonymous):

actually : (-25)*(-7) = 175 x^2-32x+175 = (x-25)(x-7)

OpenStudy (anonymous):

but it's too late anyway ...

OpenStudy (anonymous):

yeah, my fault for not starting it sooner..can't blame anything but myself XD

OpenStudy (anonymous):

I have to go, thanks for the big help. bye bye